# RAVEN: Rethinking Adversarial Video Generation with Efficient Tri-plane Networks

### v1

@@ 5 Discussion
The effectiveness of RAVEN in generating content depends greatly on the choice of the generative backend. Consequently, the same design yields varied results, influenced by the network’s capacity and dataset complexity. For instance, we’ve noted that the base model of StyleGAN2 performs reasonably well on simpler datasets, such as videos featuring static objects seen from a moving camera, a scenario that can be simulated with 3D-aware GANs like eg3D. However, it struggles significantly when dealing with more intricate datasets like UCF101. We believe that the current limitation in generation quality is primarily due to the capacity of the underlying backbone network.
In our observations, we’ve seen a consistent improvement in performance as we transitioned from the StyleGAN2 backbone to StyleGAN-XL and then to StyleGAN-T. This suggests that using a higher capacity backbone, such as a diffusion model, may be an attractive option. However, diffusion models in particular come with their own set of challenges, particularly encoding video into the triplane representation is a nontrivial task. We leave it for future exploration.
Additionally, since we employ feature flow, akin to optical flow, to model foreground and background movements separately and in a disentangled manner, we inherit some complexities of optical flow. For example, in optical flow, the lack of explicit depth information necessitates modeling an additional mask to occlude the correct part of an object when two objects converge in the image plane. This can be addressed by using four planes instead of three, with one dedicated to the three spatial dimensions and another to the time dimension.
@@
Finally, it is important to highlight that RAVEN currently does not provide direct and disentangled access to the objects within a scene it generates. This is a matter to be addressed in future research. The concept of disentanglement holds promise for the development of video generative models, and we eagerly anticipate further investigations that leverage our approach to unlock its potential.

### v2

@@ 5 Discussion
The effectiveness of RAVEN depends greatly on the generative backend. Consequently, the same design yields varied results, influenced by the network’s capacity and dataset complexity. For instance, we’ve noted that the base model of StyleGAN2 performs reasonably well on simpler datasets. However, it significantly struggles with complex datasets like UCF101 or WebVid10M-Flowers. We believe that the current limitation in generation quality is primarily due to the capacity of the underlying backbone network.
In our observations, we’ve seen a consistent improvement in performance as we transitioned from the StyleGAN2 backbone to StyleGAN-XL and then to StyleGAN-T. This suggests that using a higher-capacity backbone, such as a diffusion model, may be an attractive option. However, diffusion models in particular come with their own set of challenges, particularly encoding video into the triplane representation, which is a nontrivial task. We leave it for future exploration.
Additionally, since we employ feature flow, akin to optical flow, to model foreground and background movements separately and in a disentangled manner, we inherit some complexities of optical flow. For example, in optical flow, the lack of explicit depth information necessitates modeling an additional mask to occlude the correct part of an object when two objects converge in the image plane. This can be addressed by using four planes instead of three, and thereby operating in 3 spatial dimensions, where occlusion can be handled more naturally.
@@
Finally, it is important to highlight that RAVENcurrently does not provide direct and disentangled access to the objects within a scene it generates. This is a matter to be addressed in future research. The concept of disentanglement holds promise for video generative models, and we eagerly anticipate further investigations that leverage our approach to unlock its potential. Apart from the above, we also think a valuable direction to explore would be the in-painting and inversion capability of RAVEN.
@@ 6 Conclusion
Acnowledgements
This work was supported by the VideoPredict project, FKZ: 01IS21088.